Manufacturer of a clothes dryer has found that, when the unit price is p dollars, the revenue R ( in dollars) is
R(p)= -4p^2 + 4000p
What unit price should be established for the dryer to maximize revenue? What is the maximum revenue?
I hope you can shine some light on this problem for me! Word problems are not my strong suit! Thank you!

Answer by scott8148(6628)

the graph of this function is a downward opening parabola
___ the maximum (highest point) is at the vertex, which lies on the axis of symmetry

the general equation for the axis is ___ x = -b / (2a) ___ from f(x) = ax^2 + bx + c

in this case ___ p = -4000 / [2(-4)]